OpenStudy (anonymous):

@jim_thompson5910 can you help me with something :)

jimthompson5910 (jim_thompson5910):

what do you need?

OpenStudy (anonymous):

Hugh opened a credit card at a department store with an APR of 19.15% compounded semiannually. What is the APY on this credit card? (4 points) Select one: a. 9.58% b. 20.07% c. 25.56% d. 38.30%

jimthompson5910 (jim_thompson5910):

hint: APY = (1+r/n)^(n) - 1 http://en.wikipedia.org/wiki/Annual_percentage_yield

jimthompson5910 (jim_thompson5910):

in this case r = 0.1915 (this is the APR) n = 2 (you are compounding 2 times a year)

OpenStudy (anonymous):

im not sure how to do any of this at all

jimthompson5910 (jim_thompson5910):

APY = (1+r/n)^(n) - 1 APY = (1+0.1915/2)^(2) - 1 APY = ???

OpenStudy (anonymous):

im not sure :/

jimthompson5910 (jim_thompson5910):

From here, you use a calculator. Do you have one?

jimthompson5910 (jim_thompson5910):

if you don't have one, then copy/paste that exact expression into google and it will act like a calculator and spit out the answer

OpenStudy (anonymous):

@jim_thompson5910 what do i plug in?

jimthompson5910 (jim_thompson5910):

type (1+0.1915/2)^(2) - 1 into google and tell me what you get

OpenStudy (anonymous):

I got 0.2006680625 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

getting the same thing too

OpenStudy (anonymous):

Okay:) now what?

jimthompson5910 (jim_thompson5910):

multiply by 100 to convert to a percent, so 0.200668 = 20.066% but that rounds to 20.07%

OpenStudy (anonymous):

so add 100 to 20.07?

OpenStudy (anonymous):

i mean multiply?

jimthompson5910 (jim_thompson5910):

you multiply 0.200668 by 100 to get 20.066%

jimthompson5910 (jim_thompson5910):

then you round 20.066% to get 20.07%

jimthompson5910 (jim_thompson5910):

so the APY is 20.07%

OpenStudy (anonymous):

oh oh oh hhhhhhh okay:) cn you help with some more please?:)

jimthompson5910 (jim_thompson5910):

sure, a few more

OpenStudy (anonymous):

Laura owes $4759 on a credit card with a 20.9% interest rate compounded monthly. What is the monthly payment she should make in order to pay off this debt in 3 years, assuming she does not charge any more purchases with the card? (4 points) Select one: a. $179.05 b. $995.70 c. $132.19 d. $442.90

OpenStudy (anonymous):

i just have 6 more please pretty please then im done:)

jimthompson5910 (jim_thompson5910):

check out this calculator http://www.thecalculatorsite.com/finance/calculators/credit-card-payment-calculators.php

jimthompson5910 (jim_thompson5910):

click on "monthly repayments calculator" fill out the boxes, then calculate

OpenStudy (anonymous):

what do i put into the boxes?

jimthompson5910 (jim_thompson5910):

in the "Credit card balance", you type in the amount that's owed

OpenStudy (anonymous):

i got 179.05

jimthompson5910 (jim_thompson5910):

yep, me too

jimthompson5910 (jim_thompson5910):

optionally you can use this formula P = L((r/n)*(1 + r/n)^(n*t))/((1 + r/n)^(n*t) - 1) where L = 4759 r = 0.209 n = 12 t = 3 but that will get you the same thing

OpenStudy (anonymous):

thatsa what i thought:) ummm okay what about this one

OpenStudy (anonymous):

Explain, using complete sentences, three steps one could take to pay off their debt

jimthompson5910 (jim_thompson5910):

tell me what you think

OpenStudy (anonymous):

UM for this one I would say talk to the debt people and come up with a plan?

jimthompson5910 (jim_thompson5910):

That's one step. What about another?

OpenStudy (anonymous):

To start budgeting?

jimthompson5910 (jim_thompson5910):

good, what else?

OpenStudy (anonymous):

i cant think of another one:/

jimthompson5910 (jim_thompson5910):

another step is to pay more than the minimum

jimthompson5910 (jim_thompson5910):

if you only pay the minimum, then you'll be paying off that debt for a long time and you end up paying more over time

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Add up the balances. What do you get?

OpenStudy (anonymous):

7313.42?

jimthompson5910 (jim_thompson5910):

Now add up the credit limits. What do you get?

OpenStudy (anonymous):

17500

jimthompson5910 (jim_thompson5910):

now divide the two

OpenStudy (anonymous):

0.41790971428

jimthompson5910 (jim_thompson5910):

multiply that by 100 to get ???

OpenStudy (anonymous):

c?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

yay!Newton uses a credit card with a 18.6% APR, compounded monthly, to pay for a cruise totaling $1,920.96. He can pay $720 per month on the card. What will the total cost of this purchase be? (4 points) Select one: a. $2,458.98 b. $2,278.26 c. $1,978.02 d. $1,920.96

jimthompson5910 (jim_thompson5910):

go back here http://www.thecalculatorsite.com/finance/calculators/credit-card-payment-calculators.php

jimthompson5910 (jim_thompson5910):

now click on "repayment time" and fill out the boxes

OpenStudy (anonymous):

3?

jimthompson5910 (jim_thompson5910):

what do you mean by 3

OpenStudy (anonymous):

the answer i got was 3

jimthompson5910 (jim_thompson5910):

that's how many months it will take to pay off that debt

OpenStudy (anonymous):

yes:)

jimthompson5910 (jim_thompson5910):

look next to Total to pay back:

OpenStudy (anonymous):

so c?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Explain, using complete sentences, how you would reconcile your check register with your monthly bank statement

jimthompson5910 (jim_thompson5910):

hmm sorry, not sure what that means but I'll try to look it up

OpenStudy (anonymous):

okk

jimthompson5910 (jim_thompson5910):

I'm not sure, but it sounds like it's basically checking to see if they charged your account correctly or not

jimthompson5910 (jim_thompson5910):

and making sure they didn't overcharge or charge unfair fees

jimthompson5910 (jim_thompson5910):

and making sure things match up

OpenStudy (anonymous):

what should I say for this then?

jimthompson5910 (jim_thompson5910):

not sure tbh

jimthompson5910 (jim_thompson5910):

I couldn't really find anything. But I think it's just a case to see if things match up

OpenStudy (anonymous):

ok heres the next one

jimthompson5910 (jim_thompson5910):

what's the remaining balance in this case

OpenStudy (anonymous):

$10?

jimthompson5910 (jim_thompson5910):

remaining balance = amount you owe = ??? dollars

OpenStudy (anonymous):

it doesnt say how much patti owes

jimthompson5910 (jim_thompson5910):

how much she charges is the amount she owes

OpenStudy (anonymous):

oh! $425

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

now take 3.5% of that

OpenStudy (anonymous):

how do I do that?

jimthompson5910 (jim_thompson5910):

3.5% of 425 = 0.035*425 = ???

OpenStudy (anonymous):

14.875?

OpenStudy (anonymous):

so a?

jimthompson5910 (jim_thompson5910):

yep since $14.88 is larger than $10

OpenStudy (anonymous):

yay!Glen is comparing three investment accounts offering different rates. Account A: APR of 5.80% compounding monthly Account B: APR of 5.90% compounding quarterly Account C: APR of 5.85% compounding daily Which account will give Glen at least a 6% annual yield? (4 points) Select one: a. Account A b. Account B c. Account C d. Account B and Account C

jimthompson5910 (jim_thompson5910):

use the formula APY = (1+r/n)^(n) - 1

jimthompson5910 (jim_thompson5910):

to find the APY of each account

OpenStudy (anonymous):

What do I plug in?

jimthompson5910 (jim_thompson5910):

for instance, in account A r = 0.058 n = 12

OpenStudy (anonymous):

for that one I got 0.05956694622

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I'm getting the same as well. Repeat for the other accounts

OpenStudy (anonymous):

what do i plug in for those ones?

jimthompson5910 (jim_thompson5910):

r = interest rate (but remember to convert to decimal form)